OpenStudy (anonymous):

2|x-1|-4=-2 a. -2 b. 2 c. 11/2 d. 0

OpenStudy (anonymous):

is it d

OpenStudy (anonymous):

2x - 2 - 4= - 2 2x - 6 = - 2 +6 +6 2x = 4 ------ 2 2 x = 2 B. 2

OpenStudy (anonymous):

@madiluu ???

OpenStudy (tkhunny):

You CANNOT just discard Absolute Values. You must consider them! 2|x-1|-4=-2 Add 4 2|x-1| = +2 Divide by 2 |x-1| = 1 x-1 = 1 OR x-1 = -1 -- You must observe that |1| = 1 AND |-1| = 1 x = 2 (d) OR x = 0 (b)

OpenStudy (anonymous):

which of the two is it, x =2 (d) or x = 0 (b)

OpenStudy (tkhunny):

As the question is presented, it is BOTH.

OpenStudy (anonymous):

answers are A. -2 B. 2 C. 11/2 D. 0 there aren't 2 and 0, unless it says there's more than one possible answer?

OpenStudy (tkhunny):

I don't care what the answer choices are. Both are acceptable. I worked the problem without regard to the answer choices. If one is marked as incorrect, I would be happy to assist you in the appeal process to whatever level it needs to be pursued. Feel free to try them both in the original equation. Don't guess or assume. Prove!!!!